OpenStudy (anonymous):

Solve . Write your solution using set builder notation -2(10x-9)-17x<12(6-x)

OpenStudy (anonymous):

-(20x-18)-17x<-x+12 -36x<-12 x>1/3 but what is builder notation

OpenStudy (anonymous):

-17 x-2 (-9+10 x)<12 (6-x) = -17 x-2 (-9)-2 10 x<12 6+12 (-1) x = -17 x-9 (-1) 2-2 10 x<6 12-12 x = -17 x-18 (-1)-2 10 x<72-12 x = -17 x+18-20 x<72-12 x = 18-37 x<72-12 x x>-54/25